# Reinterpreting Landauer conductance, solving the quantum measurement problem, grand unification

### v1

@@ Abstract
In a series of recent papers we have proved rigorously that time travel is a reality and very much feasible by using quantum mechanical processes. There are plenty of indirect experimental support untill a direct experiment is conducted. The process crucially depend on the reality of a local time as well as a local partial density of states (LPDOS) that can become negative very easily in the quantum regime of mesoscopic systems. Mesoscopic systems are small enough to allow us to experimentally access the intemediate regime between the classical and quantum worlds. This LPDOS is in every sense a hidden variable in quantum mechanics that does not show up in the axiomatic framework of quantum mechanics. It can be infered through physical clocks obeying quantum dynamics and can be rigorously justified from the properties of the Hilbert space that is uniquely isomorphic to the complex plane. Therefore one can naturally guess that LPDOS will have something important to say about quantum measurement as well as the unification of classical and quantum laws. We therefore undertake the exercise to show that LPDOS can very much allow us to re-interpret the enormously successful phenomenological Landauer-Buttiker formalism for mesoscopic systems and put it on firm theoretical ground as a bridge between classical and quantum mechanics, thereby unifying them. Essentialy the local time calculated quantum mechanically can dialate exactly like the proper time of relativity and be consistent with the coordinate time of relativity. Also the measured conductance of mesoscopic samples is a deterministic quantum measurement outcome from a linear superposition of states, essentially because of LPDOS, which solves the quantum measurement problem. For this we analyze the three probe conductance formula in details and give our arguements for the general case.
@@ 1 Introduction
Mesoscopic regime is that in between the quantum and classical regime due to the small size of the samples and the low temperatures at which these systems are subjected to. Recent advances in fabrication techniques has allowed us to make high mobility two dimensional electron gas (2DEG) on which lithographic techniques can be applied to fabricate a small quantum system connected to leads that connect the quantum system to the classical world [1] and that is an essential part of these systems. These leads are labelled $\alpha$, $\beta$, $\gamma\;\;\cdots$, etc. One can make the systems so small that at low temperatures the de Broglie wavelength of the electrons become comparable to the sample size and we can look into quantum effects that are not subject to thermalization and ensamble averaging. Thus we get, what we call a general mesoscopic system that help us probe deep into the interplay of classical and quantum laws. We call it general because such a mesoscopic system becomes a true ambassador to understand the intermediate regime that lies between the quantum and classical laws. Starting from such a system, we can find variants of it to probe other problems in physics that may tend towards the classical or quantum regimes.
Motivated by the developments in fabrication of such samples, there was a flury of theoretical activity, that culminated in the celebrated Landauer-Buttiker formalism to understand these small systems. The Landauer conductance formula for example, is extremely successful in giving the conductance of such small systems, that could even be a molecule connected to leads. However, conceptual problems remain and we describe below a re-interpretation of the Landauer conductance, from the view point of our recent works [2] and mentioned in the next paragraphs.
@@
With such a local state one can obviously ask the question if it has anything to do with quantum measurements. Buttiker et al tried to obtain a direct relation between conductance (a measurable quantity) formulas for mesoscopic systems and local partial density of states (LPDOS). They could successfully do it in the semi-classical regime but it did not work for quantum regimes where $\rho_{lpd}$ can become negative. So the notion of a local-state and local time in the quantum regime remained unsolved. A local time, if it exists in quantum regime, may shed light on how general relativity and quantum mechanics can be reconcilled. These are the two issues, that we will probe in this work. Namely, how the local time reconciles quantum mechanics and relativity and how the local state help solve the quantum measurement problem. The motivation is due to the fact that we could prove [2] that negativity of $\rho_{lpd}$ is physical and corresponds to a wave-packet travelling back in time. Since a wave-packet can also carry information or signal, this raises some philosophical issues. Hence a more down to earth experimental verification will be of great help and that is also what will be given in this work.
@@ 2 Re-interpreting Landauer-Buttiker formalism
Landauer’s original arguement starts from the diagram in Fig. 1 and asserts that it represents a finite physical system where a) the reservoirs are classical and the source reservoir injects a statistically large number of electrons into the left lead that constitute a current for which b) there is a density of states (DOS) given by $\frac{1}{hv}$. Assertion b) is troublesome and will be discussed in detail. There can be no contradiction between reality and assertion a) and allows us to use the axioms of statistical meachanics. In Ref. [2], we have used the axioms of statistical mechanics and current conservation across a junction to show that the current through the system at a given incident energy $E$ is $\frac{e_{0}}{h}|t(E)|^{2}$ where $t(E)$ is the quantum mechanical transmission amplitude for the rectangular barrier as shown in Fig. 1. In this approach we do not need to invoke the idea of a DOS to derive the conductance. Instead we use the freedom of choosing a normalization constant that conserves current accross a junction with an electrostatic potential, and this normalization constant can be interpreted as a DOS that accommodate the electrons. In Eq. 13 we give the original arguements of Landauer, who on the other hand assumed that this current is carried by electrons in quantum states for which there is a well defined DOS. By understanding the simple 1D picture in Fig. 1 we can generalize to realistic mesoscopic sytems because of the decoupling of transverse modes in the leads [1]. The rectangular barrier and 1D leads can be replaced by realistic potentials for which the calculation of transmission probability will be complicated and one can proceed with this simplest case of a rectangular barrier in mind [1].
@@
Unanswered questions in Landauer’s phenomenological approach: Many different experiments show that the Landauer formalism and its diverse extensions to real systems with multiple leads, finite temperature, many body effects, non-linearities, etc, remain valid even for very small lead lengths, as we will discuss. Hence the assertion b) needs to be scrutinized. i) The parameter $E$ here is externaly applied incident energy and not eigen-energy defining an energy state. How the incident energy interacts with the energy states of the system has to be treated according to accepted rigorous theory. ii) One may also ask what are the well defined states for which we have a DOS? What exactly Landauer meant when he takes the DOS available to the electrons coming from the source to be $\frac{1}{hv}$. States of exactly which system? Are they the states of the isolated left lead or the states in the leads are the asymptotic states of a quantum mechanical scattering. Isolated left lead may be thought of as an infinite ballistic system although Landauer conductance works for fairly small leads and besides that contradicts the idea that the mechanics of the electrons in the system is a quantum mechanical scattering process. In formal scattering theory, one cannot talk of DOS for asymptotic states or DOS for non-asymptotic states. A quantum mechanical scattering has to consider the states of the full system rather than that of isolated parts. In quantum mechanics there is a well defined DOS for the phenomenon of scattering between asymptotic states that is given by the resolvent and Greens function where the exact scattering potential has to be taken into account. This is related to the dwell time to be discussed below (see Eq. 12), where the integration limits is not for the region of length $l$ but has to be over the entire region between the junctions with the reservoirs. In a quantum mechanical scattering, these are energy states. iii) What exactly is a state in mesoscopic system? Is it a momentum state for which we need translational invariance or an energy state for which we need absence of dissipation? Within the system, between the reservoirs, there is no dissipation snd so energy will be conserved, energy states may be taken as good quantum states, but that will contradict how Pauli exclusion principle is applied as will be discussed. iv) Besides, the term ”number of states” is sometimes misleading as one cannot count states in the continuum of states. This contradiction of terms is not a serious matter for large systems because of a mathematically rigorous prescription of regularization that can be applied to real systems but that prescription is not adaptable to finite mesoscopic systems.
@@
Semi-classical approach: In formal quantum mechanics only infinite systems have a continuous spectrum implying that states are uncountable. Therefore we can start from some finite system to say DOS is defined as the number of states per unit energy per unit length, which is then regularized to give measure or volume of states of the full system at an energy devided by the volume of the system to mean that at a given point or cross section. Mesoscopic systems being finite by definition, do not allow this rigorous prescription of regularization and yet electrons coming from the classical reservoir has a continuously varying incident energy. The prescription followed by S. Datta [1] based on use of Dirac delta function is ad hoc. Dirac delta function is good for solving linear differential equations but it is difficult to translate what a Dirac delta function means for finite quantum systems in nature and whether it helps us bypass the standard mathematical rules of regularization by introducing a complex self-energy parameter for a finite system. For the system size tending to infinity, there is a very close connection between the so called von Neumann’s approach through Hilbert space and Dirac’s approach through bra-ket notation. One being rigorously mathematical and the other being more intuitive, with the caution that the intuitive approach may be very helpful but can overlook errors. Dirac’s approach looses its effectiveness unless space is discritized and the Hamiltonian is represented as a matrix. Thus the self-adjoint Hamiltonian operator is reduced to a Hermitian matrix to give Heisenberg’s matrix mechanics, and then the continuum limit can be imposed through a hopping parameter without following the mathematical rigour of taking limits. If one is ready to accept some parameter fitting then this method adopted by S. Datta [1] is good for numerical calculations and hopefully that is adequate in the semi-classical regime. One may state that such ad-hoc approaches cannot reconcile the theories of quantum mechanics and relativity. One may ask what role does the hopping parameter or the self energy parameter play in general theory of relativity and how do they transform. Besides as it has been shown [2] that one cannot define a momentum oprator for a general mesoscopic system in the quantum regime, canonical quantization will not work and von Neumann’s wave-mechanics is essential.
@@
Finally, a wave-packet (also a Hilbert space element at an instant of time) is what propagates and has a time dependent equation of motion. The wavelets of the wave-packet have wave-vectors that vary continuously. While discrete states can be counted, continuum of states has a density. We can completely ignore the notion of DOS to arrive at local time recorded by a physical clock (a Larmor clock) [3]. Larmor clock devided by $\hbar$ give DOS which is completely consistent with the DOS we define quantum mechanically [3]. This DOS can be seen as how many wave-vector values are needed to make a full topological circle in a complex plane, say of the complex transmission amplitude $t$ ($Im(t)$ versus $Re(t)$) [2], as will be explained below. Thus the idea of a quantum number is replaced by winding number in this complex plane. Thus there is something like a local quantum number at the point r which is more appropriately a winding number $\delta\theta_{r}$ or $\delta\theta_{t}$ for which there is a local DOS (called local partial density of states or LPDOS)
@@
for the reflected and the transmitted electrons, respectively. Here $\theta_{r}$ and $\theta_{t}$ are the phases of the complex reflection amplitude $r$ and transmission amplitude $t$ , respectively, that change with $k$ and hence $U(\textbf{r})$, $U(\textbf{r})$ being the electro-static potential at the point r and the derivatives are functional derivatives. We have shown that this can be justified through physical clocks as well as functional analysis, without relying on operators in quantum mechanics and can be seen in more detail in [2]. At low energy $U(\textbf{r})$ is classical with no quantum corrections while only the particles are fully subject to quantum mechanics. However, in appropriate conditions, $U(\textbf{r})$ can exhibit quantum fluctuations for which there is a well established theory. This local DOS in Eq. 1 is specifically called local partial density of states (LPDOS) as it holds for only reflected electrons or only transmitted electrons. The objective reality of LPDOS can be established from functional analysis without demanding quantum states. It is consistent with quantum mechanics, but it cannot be found within the axiomatic framework of quantum mechanics [2]. The winding numbers gradually evolve into discrete quantum numbers of appropriate operators, as we pinch off the system and isolate it from external incidence from a classical reservoir.
@@
As $\theta_{t}$ is independent of a vanishingly small $U(\textbf{r})$ in the lead acting on the partial +k+k mode, that is $\frac{\delta(\theta_{t})}{\delta U(\textbf{r})}\bigg|_{\textbf{r}\in\mbox{(right lead})}=0$. This $U(\textbf{r})$ is not meant to reflect a transmitted electron but only meant to give a derivative (like a tangent to a curve does not need to increament the curve). Note that Eq. 7 is not an equality but what it says is that increasing the incident energy for a fixed $U$ in the lead, is equivalent to decreasing the $U$ in the lead for a fixed incident energy. Therefore from Eqs. 6 and 7 and noting the fact that $k$ is constant for all r in the lead ($k$ will however have a gradiant with $U(\textbf{r})$ with $E$ constant or will have a gradient with E for $U$ constant),
@@
Thus from Eq. 11, we have local partial states at a point on the right lead with a density given by $\rho_{lpd}^{t}=\frac{1}{hv}$ and for the extended region of the right lead, we get partial states with a density given by $\rho_{pd}^{t}=\frac{l}{hv}$ (from Eq. 10) and both are equally physical because LPDOS is a physical object that is coordinate dependent in general although not in the leads, that add up to give PDOS. This is thus different from DOS for the states of the full system that we get from axiomatic quantum mechanics that can be mathematically devided by the volume of the system to mean something. To be noted that, $\rho_{lpd}^{t}$ is due to local time spent at a point and $\rho_{pd}^{t}$ is due to sum of local times over the region $l$, and both comply with Lorentz’s trnsformation. At every point r and we can apply Pauli exclusion principle to the mode +k+k. So the LPDOS for +k+k mode in the right lead turns out to be the same at all points on the right lead, is $\frac{1}{hv}$, and it is subject to Pauli exclusion principle. The same is similarly true for the reflected electrons with wave-vector k-k. However, for r inside the system with a constant potential barrier, $\frac{d\theta_{t}}{dE}\not\equiv\frac{1}{e_{0}}\frac{\delta}{\delta U}\theta_{t}$ because of linear superposition of $\pm k$-modes and there is no decoupling of the $\pm k$ modes due to multiple Feynman paths. For example, $\frac{1}{e_{0}}\frac{\delta}{\delta U(\textbf{r})}\theta_{t}$ will be r dependent, depending on how close we are to the edge of the constant potential barrier. In the leads, this superposition is destroyed by the classical reservoirs [2]. For a non-rectangular barrier, say for a wavy potential barrier, inside the system there is no idnetifiable $\pm k$ states, $\frac{d\theta_{t}}{dE}\not\equiv\frac{1}{e_{0}}\frac{\delta}{\delta U}\theta_{t}$, but the leads can again be treated exactly like that of the rectangular barrier.
@@
and devided by $h$ a DOS. Here $l$ can be aribitrarily small or large. The dwell time can be found within the axiomatic framework of quantum mechanics as in a general situation (including the situation in a lead where superposition of states is destroyed by a reservoir) $\psi(k,x)$ inside the system does not have any information of from where the electrons come and where they go, but not LPDOS.
Outline: While the general view is that quantum mechanical laws are the fundamental laws of nature and classicality emerges, we adopt the view that a quantum event can only be defined between two classical events in the reservoirs and the quantum laws can coexist with classical laws without any contradiction. This has some obvious but significant consequences and solves some long standing problems in physics. First is that theory of relativity and quantum mechanics can co-exist consistently and there is no immediate need to quantize gravity, besides the fact that $U(\textbf{r})$ has well known quantum behavior. That is because now in quantum mechanics, like in theory of relativity, we can relate time intervals to signal propagation time. It was shown in ref. [2] that this signal is carried by a propagating localized undispersed quantum mechanical wave packet. We can clearly obtain a coordinate time and a proper time within the framework of quantum mechanics. The points inside the classical reservoirs have local clocks giving their respective local times or coordinate time that is consistent with mass distribution in synchronizing the classical local clocks. A signal can propagate from one reservoir to another reservoir according to the laws of quantum mechanics and the time delay defines a proper time that can dialate and contract and can be different from local times clocked by the classical clocks inside the reservoirs. The quantum system gets a space-time formulation in the sense that every point inside the quantum system has three spatial coordinates and a time coordinate. This time coordinate or coordinate time ($v\rightarrow 0$ in Eq. 3) is consistent with proper time that will be recorded by the body clock of a particle or wave-packet traversing the points and that in turn defines an un-hypheneated local quantum state that does not need a Hilbert space. Thus we get grand unification of quantum and classical laws at low energies rather than at high energies.

### v2

@@ Abstract
In a series of recent papers we have proved rigorously that time travel is a reality and very much feasible by using quantum mechanical processes. There are plenty of indirect experimental support until a direct experiment is conducted. The process crucially depend on the reality of a local time as well as a local partial density of states (LPDOS) that can become negative very easily in the quantum regime of mesoscopic systems. Mesoscopic systems are small enough to allow us to experimentally access the intermediate regime between the classical and quantum worlds. This LPDOS is in every sense a hidden variable in quantum mechanics that does not show up in the axiomatic framework of quantum mechanics. It can be inferred through physical clocks obeying quantum dynamics and can be rigorously justified from the properties of the Hilbert space that is uniquely isomorphic to the complex plane. Therefore one can naturally guess that LPDOS will have something important to say about quantum measurement as well as the unification of classical and quantum laws. We therefore undertake the exercise to show that LPDOS can very much allow us to re-interpret the enormously successful phenomenological Landauer-Buttiker formalism for mesoscopic systems and put it on firm theoretical ground as a bridge between classical and quantum mechanics, thereby unifying them. Essentially the local time calculated quantum mechanically can dilate exactly like the proper time of relativity and be consistent with the coordinate time of relativity. Also the measured conductance of mesoscopic samples is a deterministic quantum measurement outcome from a linear superposition of states, essentially because of LPDOS, which solves the quantum measurement problem. For this we analyze the three probe conductance formula in details and give our arguments for the general case.
@@ 1 Introduction
Mesoscopic regime is that in between the quantum and classical regimes due to the small size of the samples and the low temperatures at which these systems are subjected to. Recent advances in fabrication techniques has allowed us to make high mobility two dimensional electron gas (2DEG) on which lithographic techniques can be applied to fabricate a small quantum system connected to leads that connect the quantum system to the classical world [1] and that is an essential part of these systems. Supposing these leads are labelled $\alpha$, $\beta$, $\gamma\;\;\cdots$, etc. One can make the systems so small that at low temperatures the de Broglie wavelength of the electrons become comparable to the sample size and we can look into quantum effects that are not subject to thermalization and ensemble averaging. Thus we get, what we call a general mesoscopic system that help us probe deep into the interplay of classical and quantum laws. We call it general because such a mesoscopic system becomes a true ambassador to understand the intermediate regime that lies between the quantum and classical laws. Starting from such a system, we can find variants of it to probe other problems in physics that may tend towards the classical or quantum regimes.
Motivated by the developments in fabrication of such samples, there was a flurry of theoretical activity, that culminated in the celebrated Landauer-Buttiker formalism to understand these small systems. The Landauer conductance formula for example, is extremely successful in giving the conductance of such small systems, that could even be a molecule connected to leads. However, conceptual problems remain and we describe below a re-interpretation of the Landauer conductance, from the view point of our recent works [2] and mentioned in the next paragraphs.
@@
Once we know $\rho_{lpd}$ we can use it to find other relevant quantities like injectance that we find in axiomatic framework of quantum mechanics and so we are not loosing anything but only gaining some extra objects by this assertion. At any point we can go to the closed system by making the leads $\gamma$ and $\alpha$ very weakly connected to the system and recover all that we get by solving the Schrodinger equation for the closed system. Besides it may not be always necessary to pinch off the leads physically as von Neumann quantum mechanics give us a method to infer about the closed system from the scattering states as has been worked out in chapter 4 of Ref [2].
With such a local state one can obviously ask the question if it has anything to do with quantum measurements. Buttiker et al tried to obtain a direct relation between conductance (a measurable quantity) formulas for mesoscopic systems and local partial density of states (LPDOS). They could successfully do it in the semi-classical regime but it did not work for quantum regimes where $\rho_{lpd}$ can become negative. So the notion of a local-state and local time in the quantum regime remained unsolved. A local time, if it exists in quantum regime, may shed light on how general relativity and quantum mechanics can be reconciled. These are the two issues, that we will probe in this work. Namely, how the local time reconciles quantum mechanics and relativity and how the local state help solve the quantum measurement problem. The motivation is due to the fact that we could prove [2] that negativity of $\rho_{lpd}$ is physical and corresponds to a wave-packet traveling back in time. Since a wave-packet can also carry information or signal, this raises some philosophical issues. Hence a more down to earth experimental verification will be of great help and that is also what will be given in this work.
@@ 2 Re-interpreting Landauer-Buttiker formalism
Landauer’s original argument starts from the diagram in Fig. 1 and asserts that it represents a finite physical system where a) the reservoirs are classical and the source reservoir injects a statistically large number of electrons into the left lead that constitute a current for which b) there is a density of states (DOS) given by $\frac{1}{hv}$. Assertion b) is troublesome and will be discussed in detail. There can be no contradiction between reality and assertion a) and allows us to use the axioms of statistical meachanics. In Ref. [2], we have used the axioms of statistical mechanics and current conservation across a junction to show that the current through the system at a given incident energy $E$ is $\frac{e_{0}}{h}|t(E)|^{2}$ where $t(E)$ is the quantum mechanical transmission amplitude for the rectangular barrier as shown in Fig. 1. In this approach we do not need to invoke the idea of a DOS to derive the conductance. Instead we use the freedom of choosing a normalization constant that conserves current across a junction with an electrostatic potential, and this normalization constant can be interpreted as a DOS that accommodate the electrons. In Eq. 13 we give the original arguments of Landauer, who on the other hand assumed that this current is carried by electrons in quantum states for which there is a well defined DOS. By understanding the simple 1D picture in Fig. 1 we can generalize to realistic mesoscopic systems because of the decoupling of transverse modes in the leads [1]. The rectangular barrier and 1D leads can be replaced by realistic potentials for which the calculation of transmission probability will be complicated and one can proceed with this simplest case of a rectangular barrier in mind [1].
@@
Unanswered questions in Landauer’s phenomenological approach: Many different experiments show that the Landauer formalism and its diverse extensions to real systems with multiple leads, finite temperature, many body effects, non-linearities, etc, remain valid even for very small lead lengths, as we will discuss. Hence the assertion b) needs to be scrutinized. i) The parameter $E$ here is externally applied incident energy and not eigen-energy defining an energy state. How the incident energy interacts with the energy states of the system has to be treated according to accepted rigorous theory. ii) One may also ask what are the well defined states for which we have a DOS? What exactly Landauer meant when he takes the DOS available to the electrons coming from the source to be $\frac{1}{hv}$. States of exactly which system? Are they the states of the isolated left lead or the states in the leads are the asymptotic states of a quantum mechanical scattering. Isolated left lead may be thought of as an infinite ballistic system although Landauer conductance works for fairly small leads and besides that contradicts the idea that the mechanics of the electrons in the system is a quantum mechanical scattering process. In formal scattering theory, one cannot talk of DOS for asymptotic states or DOS for non-asymptotic states. A quantum mechanical scattering has to consider the states of the full system rather than that of isolated parts. In quantum mechanics there is a well defined DOS for the phenomenon of scattering between asymptotic states that is given by the resolvent and Greens function where the exact scattering potential has to be taken into account. This is related to the dwell time to be discussed below (see Eq. 12), where the integration limits is not for the region of length $l$ but has to be over the entire region between the junctions with the reservoirs. In a quantum mechanical scattering, these are energy states. iii) What exactly is a state in mesoscopic system? Is it a momentum state for which we need translational invariance or an energy state for which we need absence of dissipation? Within the system, between the reservoirs, there is no dissipation and so energy will be conserved, energy states may be taken as good quantum states, but that will contradict how Pauli exclusion principle is applied as will be discussed. iv) Besides, the term ”number of states” is sometimes misleading as one cannot count states in the continuum of states. This contradiction of terms is not a serious matter for large systems because of a mathematically rigorous prescription of regularization that can be applied to real systems but that prescription is not adaptable to finite mesoscopic systems.
@@
Semi-classical approach: In formal quantum mechanics only infinite systems have a continuous spectrum implying that states are uncountable. Therefore we can start from some finite system to say DOS is defined as the number of states per unit energy per unit length, which is then regularized to give measure or volume of states of the full system at an energy divided by the volume of the system to mean that at a given point or cross section. Mesoscopic systems being finite by definition, do not allow this rigorous prescription of regularization and yet electrons coming from the classical reservoir has a continuously varying incident energy. The prescription followed by S. Datta [1] based on use of Dirac delta function is ad hoc. Dirac delta function is good for solving linear differential equations but it is difficult to translate what a Dirac delta function means for finite quantum systems in nature and whether it helps us bypass the standard mathematical rules of regularization by introducing a complex self-energy parameter for a finite system. For the system size tending to infinity, there is a very close connection between the so called von Neumann’s approach through Hilbert space and Dirac’s approach through bra-ket notation. One being rigorously mathematical and the other being more intuitive, with the caution that the intuitive approach may be very helpful but can overlook errors. Dirac’s approach looses its effectiveness unless space is discritized and the Hamiltonian is represented as a matrix. Thus the self-adjoint Hamiltonian operator is reduced to a Hermitian matrix to give Heisenberg’s matrix mechanics, and then the continuum limit can be imposed through a hopping parameter without following the mathematical rigor of taking limits. If one is ready to accept some parameter fitting then this method adopted by S. Datta [1] is good for numerical calculations and hopefully that is adequate in the semi-classical regime. One may state that such ad-hoc approaches cannot reconcile the theories of quantum mechanics and relativity. One may ask what role does the hopping parameter or the self energy parameter play in general theory of relativity and how do they transform. Besides as it has been shown [2] that one cannot define a momentum operator for a general mesoscopic system in the quantum regime, canonical quantization will not work and von Neumann’s wave-mechanics is essential.
@@
Finally, a wave-packet (also a Hilbert space element at an instant of time) is what propagates and has a time dependent equation of motion. The wavelets of the wave-packet have wave-vectors that vary continuously. While discrete states can be counted, continuum of states has a density. We can completely ignore the notion of DOS to arrive at local time recorded by a physical clock (a Larmor clock) [3]. Larmor clock divided by $\hbar$ give DOS which is completely consistent with the DOS we define quantum mechanically [3]. This DOS can be seen as how many wave-vector values are needed to make a full topological circle in a complex plane, say of the complex transmission amplitude $t$ ($Im(t)$ versus $Re(t)$) [2], as will be explained below. Thus the idea of a quantum number is replaced by winding number in this complex plane. Thus there is something like a local quantum number at the point r which is more appropriately a winding number $\delta\theta_{r}$ or $\delta\theta_{t}$ for which there is a local DOS (called local partial density of states or LPDOS)
@@
for the reflected and the transmitted electrons, respectively. Here $\theta_{r}$ and $\theta_{t}$ are the phases of the complex reflection amplitude $r$ and transmission amplitude $t$ , respectively, that change with $k$ and hence $U(\textbf{r})$, $U(\textbf{r})$ being the electro-static potential at the point r and the derivatives are functional derivatives. Here $e_{0}$ is the charge of the particles. We have shown that this can be justified through physical clocks as well as functional analysis, without relying on operators in quantum mechanics and can be seen in more detail in [2]. At low energy $U(\textbf{r})$ is classical with no quantum corrections while only the particles are fully subject to quantum mechanics. However, in appropriate conditions, $U(\textbf{r})$ can exhibit quantum fluctuations for which there is a well established theory. This local DOS in Eq. 1 is specifically called local partial density of states (LPDOS) as it holds for only reflected electrons or only transmitted electrons. The objective reality of LPDOS can be established from functional analysis without demanding quantum states. It is consistent with quantum mechanics, but it cannot be found within the axiomatic framework of quantum mechanics [2]. The winding numbers gradually evolve into discrete quantum numbers of appropriate operators, as we pinch off the system and isolate it from external incidence from a classical reservoir.
@@
As $\theta_{t}$ is independent of a vanishingly small $U(\textbf{r})$ in the lead acting on the partial +k+k mode, that is $\frac{\delta(\theta_{t})}{\delta U(\textbf{r})}\bigg|_{\textbf{r}\in\mbox{(right lead})}=0$. This $U(\textbf{r})$ is not meant to reflect a transmitted electron but only meant to give a derivative (like a tangent to a curve does not need to increment the curve). Note that Eq. 7 is not an equality but what it says is that increasing the incident energy for a fixed $U$ in the lead, is equivalent to decreasing the $U$ in the lead for a fixed incident energy. Therefore from Eqs. 6 and 7 and noting the fact that $k$ is constant for all r in the lead ($k$ will however have a gradient with $U(\textbf{r})$ with $E$ constant or will have a gradient with E for $U$ constant),
@@
Thus from Eq. 11, we have local partial states at a point on the right lead with a density given by $\rho_{lpd}^{t}=\frac{1}{hv}$ and for the extended region of the right lead, we get partial states with a density given by $\rho_{pd}^{t}=\frac{l}{hv}$ (from Eq. 10) and both are equally physical because LPDOS is a physical object that is coordinate dependent in general although not in the leads, that add up to give PDOS. This is thus different from DOS for the states of the full system that we get from axiomatic quantum mechanics that can be mathematically divided by the volume of the system to mean something. To be noted that, $\rho_{lpd}^{t}$ is due to local time spent at a point and $\rho_{pd}^{t}$ is due to sum of local times over the region $l$, and both comply with Lorentz’s trnsformation. At every point r and we can apply Pauli exclusion principle to the mode +k+k. So the LPDOS for +k+k mode in the right lead turns out to be the same at all points on the right lead, is $\frac{1}{hv}$, and it is subject to Pauli exclusion principle. The same is similarly true for the reflected electrons with wave-vector k-k. However, for r inside the system with a constant potential barrier, $\frac{d\theta_{t}}{dE}\not\equiv\frac{1}{e_{0}}\frac{\delta}{\delta U}\theta_{t}$ because of linear superposition of $\pm k$-modes and there is no decoupling of the $\pm k$ modes due to multiple Feynman paths. For example, $\frac{1}{e_{0}}\frac{\delta}{\delta U(\textbf{r})}\theta_{t}$ will be r dependent, depending on how close we are to the edge of the constant potential barrier. In the leads, this superposition is destroyed by the classical reservoirs [2]. For a non-rectangular barrier, say for a wavy potential barrier, inside the system there is no idnetifiable $\pm k$ states, $\frac{d\theta_{t}}{dE}\not\equiv\frac{1}{e_{0}}\frac{\delta}{\delta U}\theta_{t}$, but the leads can again be treated exactly like that of the rectangular barrier.
@@
and divided by $h$ a DOS. Here $l$ can be aribitrarily small or large. The dwell time can be found within the axiomatic framework of quantum mechanics as in a general situation (including the situation in a lead where superposition of states is destroyed by a reservoir) $\psi(k,x)$ inside the system does not have any information of from where the electrons come and where they go, but not LPDOS.
Outline: While the general view is that quantum mechanical laws are the fundamental laws of nature and classicality emerges, we adopt the view that a quantum event can only be defined between two classical events in the reservoirs and the quantum laws can coexist with classical laws without any contradiction. This has some obvious but significant consequences and solves some long standing problems in physics. First is that theory of relativity and quantum mechanics can co-exist consistently and there is no immediate need to quantize gravity, besides the fact that $U(\textbf{r})$ has well known quantum behavior. That is because now in quantum mechanics, like in theory of relativity, we can relate time intervals to signal propagation time. It was shown in ref. [2] that this signal is carried by a propagating localized undispersed quantum mechanical wave packet. We can clearly obtain a coordinate time and a proper time within the framework of quantum mechanics. The points inside the classical reservoirs have local clocks giving their respective local times or coordinate time that is consistent with mass distribution in synchronizing the classical local clocks. A signal can propagate from one reservoir to another reservoir according to the laws of quantum mechanics and the time delay defines a proper time that can dilate and contract and can be different from local times clocked by the classical clocks inside the reservoirs. The quantum system gets a space-time formulation in the sense that every point inside the quantum system has three spatial coordinates and a time coordinate. This time coordinate or coordinate time ($v\rightarrow 0$ in Eq. 3) is consistent with proper time that will be recorded by the body clock of a particle or wave-packet traversing the points and that in turn defines an un-hypheneated local quantum state that does not need a Hilbert space. Thus we get grand unification of quantum and classical laws at low energies rather than at high energies.